### v1

# Per-Gaussian Embedding-Based Deformation for Deformable 3D Gaussian Splatting

Abstract
As 3D Gaussian Splatting (3DGS) provides fast and high-quality novel view synthesis, it is a natural extension to deform a canonical 3DGS to multiple frames. However, previous works fail to accurately reconstruct dynamic scenes, especially 1) static parts moving along nearby dynamic parts, and 2) some dynamic areas are blurry. We attribute the failure to the wrong design of the deformation field, which is built as a coordinate-based function. This approach is problematic because 3DGS is a mixture of multiple fields centered at the Gaussians, not just a single coordinate-based framework. To resolve this problem, we define the deformation as a function of per-Gaussian embeddings and temporal embeddings. Moreover, we decompose deformations as coarse and fine deformations to model slow and fast movements, respectively. Also, we introduce an efficient training strategy for faster convergence and higher quality. Project page: https://jeongminb.github.io/e-d3dgs/
Keywords:
Gaussian splatting Dynamic scene reconstruction Novel view synthesis1 Introduction
Dynamic scene reconstruction from multi-view input videos is an important task in computer vision, as it can be extended to various applications and industries such as mixed reality, content production, etc. Neural Radiance Fields (NeRF) [14], which enable photorealistic novel view synthesis from multi-view inputs, can represent dynamic scenes by modeling the scene with an additional time input [19, 10]. However, typical NeRFs require querying multilayer perceptron (MLP) for hundreds of points per camera ray, which limits rendering speed.
On the other hand, the recently emerging 3D Gaussian Splatting (3DGS) [8] has the advantage of real-time rendering compared to NeRFs using a differentiable rasterizer for 3D Gaussian primitives. 3DGS directly optimizes the parameters of 3D Gaussian (position, opacity, anisotropic covariance, and spherical harmonics (SH) coefficients) and renders them via projection and -blending. Since 3DGS has the characteristics of continuous volumetric radiance fields, some recent studies[25, 28, 3, 7, 12, 29] represent dynamic scenes by defining a canonical 3DGS and deforming it to individual frames as deformable NeRFs [28] do. Specifically, they model the deformation as a function of 4D (x, y, z, t) coordinates with MLPs or grids, i.e., fields, predicting the change in the 3D Gaussian parameters.
However, it is not appropriate to model the deformation of Gaussian parameters as a field for representing dynamic scenes because the positions of Gaussians change during optimization. Furthermore, nearby Gaussians tend to have similar deformations because the deformation function is piecewise continuous. As shown in Figure 1, the Gaussians in static regions, e.g., window, wall, and handle, are attracted by the motion of nearby dynamic objects. Moreover, the existing deformation-based approaches have an inherent property: all frames share a single canonical set of Gaussians. It makes the deformation hard to reconstruct different types or difficulties of movements in the dynamic scenes.
In this paper, we model the deformation of Gaussians at frames as 1) a function of a product space of per-Gaussian embeddings and temporal embeddings. We expect this rational design to bring quality improvement by precisely modeling different deformations of different Gaussians without interference due to adjacency. Additionally, 2) We decompose temporal variations of the parameters into coarse and fine components, namely coarse-fine deformation. The coarse deformation represents large or slow movements in the scene, while fine deformation learns the fast or detailed movements that coarse deformation does not cover. Finally, we propose 3) an efficient training strategy that samples difficult parts of the scene more frequently and facilitates the densification of Gaussians.
In our experiments, we observe that our per-Gaussian embeddings and coarse-fine deformation improve the deformation quality, while an efficient training strategy enables faster convergence. Additionally, our approach outperforms baselines in representing dynamic scenes, capturing fine details in dynamic regions, and excelling even under challenging camera settings.
2 Related Work
In this section, we review methods for dynamic scene reconstruction that deforms 3D canonical space and methods for reconstructing dynamic scenes utilizing dynamic 3D Gaussians. Afterward, we review the methods using embedding and efficient sampling techniques.
2.0.1 Deforming 3D Canonical Space
D-NeRF[19] reconstructs dynamic scenes by deforming ray samples over time, using the deformation network that takes 3D coordinates and timestamps of the sample as inputs. Nerfies[16] and HyperNeRF[17] use per-frame trainable deformation codes instead of time conditions to deform the canonical space. Instead of deforming from the canonical frame to the entire frames, HyperReel[1] deforms the ray sample of the keyframe to represent the intermediate frame. 4DGaussians [25] and D3DGS [28] reconstruct the dynamic scene with a deformation network which inputs the center position of the canonical 3D Gaussians and timestamps. In contrast, we demonstrate a novel deformation scheme to represent the deformation as a function of a product space of per-Gaussian latent embeddings and temporal embeddings.
2.0.2 Dynamic 3D Gaussians
To extend the fast rendering speed of 3D Gaussian Splatting[8] into dynamic scene reconstructions. 4DGaussians[25] decodes features from multi-resolution HexPlanes for temporal deformation of 3D Gaussians. While D3DGS[28] uses an implicit function that processes the time and location of the Gaussian. 4DGS[26] decomposed the 4D Gaussians into a time-conditioned 3D Gaussians and a marginal 1D Gaussians. STG[11] represents changes in 3D Gaussian over time through a temporal opacity and a polynomial function for each Gaussian.
2.0.3 Latent Embedding on Novel View Synthesis
Some studies incorporate latent embeddings to represent different states of the static and dynamic scene. NeRF-W[13] and Block-NeRF[22] employ per-image appearance embeddings to capture different states of a scene, representing the dynamic scenes from in-the-wild image collections. DyNeRF and MixVoxels[10, 23] employ a temporal embedding for each frame to represent dynamic scenes. Nerfies[16] and HyperNeRF[17] incorporate both per-frame appearance and deformation embeddings. Sync-NeRF[9] introduces time offset to calibrate the misaligned temporal embeddings on dynamic scenes from unsynchronized videos. We introduce per-Gaussian latent embedding to encode the changes over time of each Gaussian, and use temporal embeddings to represent different states in each frame of the scene.
2.0.4 Efficient Training Strategy
To accelerate convergence and enhance performance, ActiveNeRF[15] selects the view by reducing the uncertainty of the input views. DS-NeRF and TermiNeRF[2, 18] samples the points with the sampling network. In the dynamic scene reconstruction, DyNeRF[10] samples rays over time, calculating he weight of each ray based on the residual difference between its color across time. STG[11] samples additional Gaussians along the ray within patches exhibiting significant errors. On the other hand, we present a simple view sampling method through pairwise camera distances, for the first time, we present a sampling method for a target training frame.
3 Method
In this section, we first provide a brief overview of 3D Gaussian Splatting (Section 3.1). Next, we introduce our overall framework, embedding-based deformation for Gaussians (Section 3.2) and coarse-fine deformation scheme consisting of coarse and fine deformation functions (Section 3.3). Finally, we present an efficient training strategy for faster convergence and higher performance (Section 3.4).
3.1 Preliminary: 3D Gaussian Splatting
3D Gaussian splatting[8] optimizes a set of anisotropic 3D Gaussians through differentiable tile rasterization to reconstruct a static 3D scene. By its efficient rasterization, the optimized model enables real-time rendering of high-quality images. Each 3D Gaussian function at the point consist with position , rotation , and scale :.
To projecting 3D Gaussians to 2D for rendering, covariance matrix are calculated by viewing transform and the Jacobian of the affine approximation of the projective transfomation[31] as follows:
Blending depth-ordered projected points that overlap the pixel, the Gaussian function is multiplied by the opacity of the Gaussian and calculates the pixel color with the color of the Gaussian :
The color of Gaussian is determined using the SH coefficient with view direction.
3.2 Embedding-Based Deformation for Gaussians
Deformable NeRFs consist of a deformation field that predicts displacement for a given coordinate from the canonical space to each target frame, and a radiance field that maps color and density from a given coordinate in the canonical space (). Existing deformable Gaussian methods employ the same approach for predicting the deformation of Gaussians, i.e., utilizing a deformation field based on coordinates.
Unlike previous methods, we start from the design of 3DGS: the 3D scene is represented as a mixture of Gaussians that have individual radiance fields. Accordingly, the deformation should be defined for each Gaussian. Based on this intuition, we introduce a function that produces deformation from learnable embeddings belonging to individual Gaussians (Figure 2a), and typical temporal embeddings for different frames:
where is a rotation quaternion, is a vector for scaling, is an opacity, and is SH coefficients for modeling view-dependent color. We implement as a shallow multi-layer perceptron (MLP) followed by an MLP head for each parameter. As a result, the Gaussian parameters at frame are determined by adding to the canonical Gaussian parameters (Figure 2c).
We jointly optimize the per-Gaussian embeddings , the deformation function , and the canonical Gaussian parameters to minimize rendering loss. We use the L1 and periodic DSSIM as the rendering loss between the rendered image and the ground truth image.
3.3 Coarse-Fine Deformation
Different parts of a scene may have coarse and fine motions [5]. E.g., a hand swiftly stirs a pan (fine) while a body slowly moves from left to right (coarse). Based on this intuition, we introduce a coarse-fine deformation that produces a summation of coarse and fine deformations.
Coarse-fine deformation consists of two functions with the same architecture: one for coarse and one for fine deformation (Figure 2c). The functions receive different temporal embeddings as follows:
Following typical temporal embeddings, we start from a 1D feature grid for frames and use an embedding for fine deformation. For coarse deformation, we linearly downsample by a factor of 5 to remove high-frequencies responsible for fast and detailed deformation. Then we compute as a linear interpolation of embeddings at enclosing grid points (Figure 2b).
As a result, coarse deformation (, ) is responsible for representing large or slow movements in the scene, while fine deformation (, ) learns the fast or detailed movements that coarse deformation does not cover. This improves the deformation quality. Refer to the Ablation study section for more details.
3.4 Efficient Training Strategy
Last but not least, we introduce a training strategy for faster convergence and higher performance. The first is to evenly cover multi-view camera perspectives and exclude the camera index that was used in the previous iteration. We pre-compute pairwise distances of all camera origins before the start of the training. Then we exclude the cameras with distance to the previous camera less than 40 percentile of all distances.
The second is to sample frames within the target cameras. Different frames may have different difficulties for reconstruction. We measure the difficulty of a frame by the magnitude of error at the frame. Then, we sample training frames with a categorical distribution which is proportional to the magnitude of error. For multi-view videos, we average the errors from all viewpoints. We use this sampling every other iteration with random sampling.
The third is to change the policy for Gaussian densification. Previous methods invoke the densification by minimizing L1 loss [25] or L1 loss and DSSIM loss for every iteration [27, 28, 11]. We observe that the DSSIM loss improves the visual quality in the background but takes longer training time. Therefore, we use L1 loss for every iteration and periodically use the additional multi-view DSSIM loss in the frame with high error. This encourages densification in the region where the model struggles (See Section 0.A.1).
4 Experiment
In this section, we first describe the criterion for selection of baselines, and evaluation metrics. We then demonstrate the effectiveness of our method through comparisons with various baselines and datasets (Section 4.1-4.2). Finally, we conduct analysis and ablations of our method (Section 4.3).
4.0.1 Baselines
We choose the start-of-the-art method as a baseline in each datasets. We compared against DyNeRF, NeRFPlayer, MixVoxels, K-Planes, HyperReel, Nerfies, HyperNeRF, and TiNeuVox on the NeRF baseline. In detail, we use the version of NeRFPlayer TensoRF VM, HyperNeRF DF, Mixvoxels-L, K-Planes hybrid. We compared with 4DGaussians, 4DGS, and D3DGS based on the Gaussian baseline. Meanwhile, we have not included STG in our comparison due to its requirement for per-frame Structure from Motion (SfM) points, which makes conducting a fair comparison challenging. Also, STG is not a deformable 3D Gaussian approach. We followed the official code and configuration, except increasing the training iterations for the Technicolor dataset to 1.5 times that of the 4DGaussians, in comparison to the Neural 3D Video dataset.
4.0.2 Metrics
We report the quality of rendered images using PSNR, SSIM, and LPIPS. Peak Signal-to-Noise Ratio (PSNR) quantifies pixel color error between the rendered video and the ground truth. We utilize SSIM [24] to account for the perceived similarity of the rendered image. Additionally, we measure higher-level perceptual similarity using Learned Perceptual Image Patch Similarity (LPIPS) [30] with an AlexNet Backbone. Higher PSNR and SSIM values and lower LPIPS values indicate better visual quality.
4.1 Effectiveness on dynamic region
4.1.1 Neural 3D Video Dataset
[10] includes 20 multi-view videos, with each scene consisting of either 300 frames, except for the flame_salmon scene, which comprises 1200 frames. These scenes encompass a relatively long duration and various movements, with some featuring multiple objects in motion. We utilized Neural 3D Video dataset to observe the capability to capture dynamic areas. Total six scenes (coffee_martini, cook_spinach, cut_roasted_beef, flame_salmon, flame_steak, sear_steak) are evaluated in Figure 4 and Table 1. The flame_salmon scene divided into four segments, each containing 300 frames.
Figure 4 reports the rendering quality. Our method successfully reconstructs the fine details in moving areas, outperforming baselines on average metrics across test views. Baselines show blurred dynamic areas or severe artifacts in low-light scenes such as cook_spinach and flame_steak. 4DGS exhibits the disappearance of some static areas. In 4DGaussians, a consistent over-smoothing occurs in dynamic areas. All baselines experienced reduced quality in reflective or thin dynamic areas like clamps, torches, and windows.
Table 1 presents quantitative metrics on the average metrics across the test views of all scenes, computational and storage costs on the first fragment of flame_salmon scene. Refer to the Appendix for per-scene details. Our method demonstrates superior reconstruction quality across all metrics compared to baselines. As the table shows, NeRF baselines generally required longer training and rendering times. While 4DGS shows relatively high reconstruction performance, it demands longer training times and larger VRAM storage capacity compared to other baselines. 4DGaussians requires lower computational and storage cost but it displays low reconstruction quality in some scenes with rapid dynamics, as shown in the teaser and Figure 4.
| model | metric | computational cost | ||||
| PSNR | SSIM | LPIPS | Training time | FPS | Model size | |
| DyNeRF12 [10] | 29.58 | - | 0.083 | 1344 hours | 0.01 | 56 MB |
| NeRFPlayer13 [21] | 30.69 | - | 0.111 | 6 hours | 0.05 | 1654 MB |
| MixVoxels [23] | 30.30 | 0.918 | 0.127 | 1 hours 40 mins | 0.93 | 512 MB |
| K-Planes [6] | 30.86 | 0.939 | 0.096 | 1 hours 40 mins | 0.13 | 309 MB |
| HyperReel4 [1] | 30.37 | 0.921 | 0.106 | 9 hours 20 mins | 1.0 | 1362 MB |
| 4DGS [26] | 31.19 | 0.940 | 0.051 | 9 hours 30 mins | 33.7 | 8700 MB |
| 4DGaussians [25] | 30.71 | 0.935 | 0.056 | 50 mins | 51.9 | 59 MB |
| Ours | 31.42 | 0.945 | 0.037 | 3 hours 35 mins | 19.9 | 137 MB |
4.1.2 Technicolor Light Field Dataset
[20] is a multi-view dataset captured with a time-synchronized camera rig, containing intricate details. We train ours and the baselines on 50 frames of five commonly used scenes (Birthday, Fabien, Painter, Theater, Trains) using full-resolution videos at pixels, with the second row and second column cameras used as test views.
As shown in Figure 5, HyperReel produces noisy artifacts due to incorrect predictions of the displacement vector. 4DGaussians fails to capture fine details in dynamic areas, exhibiting over-smoothing results. All baselines struggle to accurately reconstruct rapidly moving thin areas like fingers.
Table 2 reports the average metrics across the test views of all scenes, computational and storage costs on Painter scene. HyperReel demonstrates overall high-quality results but struggles with relatively slow training times and FPS, and a larger model size. 4DGaussians exhibits fast training times and FPS but significantly underperforms in reconstructing fine details compared to other baselines. However, our method demonstrates superior reconstruction quality and faster FPS compared to the baselines.
| model | metric | computational cost | ||||
| PSNR | SSIM | LPIPS | Training time | FPS | Model size | |
| DyNeRF | 31.80 | - | 0.140 | - | 0.02 | 0.6 MB |
| HyperReel | 32.32 | 0.899 | 0.118 | 2 hours 45 mins | 0.91 | 289 MB |
| 4DGaussians | 29.62 | 0.844 | 0.176 | 25 mins | 34.8 | 51 MB |
| Ours | 33.38 | 0.907 | 0.103 | 1 hours 30 mins | 36.5 | 56 MB |
4.2 Scenes with Challenging Camera Setting
4.2.1 HyperNeRF dataset
includes videos using two Pixel 3 phones rigidly mounted on a handheld capture rig. We train on all frames of four scenes (3D Printer, Banana, Broom, Chicken) at a resolution downsampled by a factor of two to . Due to memory constraints, D3DGS is trained on images further downsampled to a quarter of their original size.
Figure 6 shows that the previous methods struggle to reconstruct fast-moving parts such as fingers and broom. Especially D3DGS deteriorates in Broom. Table 3 reports the average metrics across the test views of all scenes, computational and storage costs on Broom scene. The table shows that our method outperforms the reconstruction performance with previous methods along with compact model size and faster FPS.
| model | metric | computational cost | ||||
| PSNR | SSIM | LPIPS | Training time | FPS | Model size | |
| Nerfies [16] | 22.23 | - | 0.170 | hours | 1 | - |
| HyperNeRF DS [17] | 22.29 | 0.598 | 0.153 | 32 hours | 1 | 15 MB |
| TiNeuVox [4] | 24.20 | 0.616 | 0.393 | 3 hours 30 mins | 1 | 48 MB |
| D3DGS [28] | 22.40 | 0.598 | 0.275 | 16 mins | 6.95 | 309 MB |
| 4DGaussians | 25.03 | 0.682 | 0.281 | 50 mins | 96.3 | 60 MB |
| Ours | 25.53 | 0.694 | 0.242 | 1 hours 20 mins | 30.9 | 34 MB |
4.3 Analyses and Ablation Study
4.3.1 Improving Rendering Speed
Our method can enhance rendering speed by omitting the deformation of Gaussians in static areas, which only require a single computation of parameter changes from the canonical space. To identify static Gaussians, we sum up the L1 norms of each Gaussian parameter over time and select Gaussians below the percentile based on the aggregated value. As Table 4 reports, we can obtain almost identical rendering quality to the original, despite omitting decoder evaluations for 90% of the total Gaussian. The experiments are conducted on flame_salmon_1 scene with a single NVIDIA RTX 4090.
| Skip deformation | PSNR | SSIM | LPIPS | FPS |
| 0% (original) | 29.72 | 0.935 | 0.038 | 17.1 |
| 50% | 29.72 | 0.935 | 0.038 | 25.7 |
| 70% | 29.72 | 0.935 | 0.038 | 26.0 |
| 90% | 29.68 | 0.935 | 0.038 | 28.4 |
4.3.2 Deformation components
In Figure 7, we present an analysis of the coarse-fine deformation. To achieve this, we render a flame_steak scene by omitting each of our deformation components one by one. Our full rendering results from adding coarse and fine deformation to the canonical space (Figure 7d). When both are removed, rendering yields an image in canonical space (Figure 7a). Rendering with the coarse deformation, which handles large or slow changes in the scene, produces results similar to the full rendering (Figure 7b). On the other hand, fine deformation is responsible for fast or detailed changes in the scene, yielding rendering similar to canonical space (Figure 7c).
To examine the roles of the coarse and fine deformation in the coarse-fine deformation, we conduct a visualization on flame_steak scene. First, we compute the Euclidean norm of positional shifts between the current and subsequent frames. We then add the value to the DC components of the SH coefficients proportionally to the magnitude: blue for coarse deformation and red for fine deformation. For visual clarity, we render the original scene in grayscale. As illustrated in Figure 8, coarse deformation models slower changes such as body movement, while fine deformation models faster movements like cooking arms. Thus, we demonstrate that by downsampling the temporal embedding grid , we can effectively separate and model slow and fast deformations in the scene.
| Method | PSNR | SSIM | LPIPS |
| Ours Full | 29.08 | 0.924 | 0.056 |
| w/o Coarse-Fine deformation | 28.74 | 0.922 | 0.060 |
| w/o Efficient training strategy | 28.89 | 0.920 | 0.056 |
| Method | PSNR | SSIM | LPIPS |
| Ours Full | 30.28 | 0.934 | 0.047 |
| coarse deformation only | 29.71 | 0.930 | 0.050 |
| fine deformation only | 29.62 | 0.929 | 0.048 |
| Ours + injection | 28.17 | 0.916 | 0.057 |
4.3.3 Ablation study
To verify the effectiveness of each component of our method, we report the performance when each component is removed. To observe the impact of the coarse-fine deformation, we train a total of 100K iterations on flame_salmon_4 scene. To show the efficiency of our training strategy, we evaluate the rendering results from checkpoints of 30K iterations. As shown in Table 5, we confirm that the performance decreases when each component is removed.
Furthermore, we report the results of an ablation study on the deformation decoder in Figure 9 and Table 6. First, our full method (using both coarse and fine decoders) produces clear rendering results and models dynamic variations well (Figure 9a). Training only with the coarse decoder leads to blurred dynamic areas and a failure to accurately capture detailed motion (Figure 9)b. On the other hand, only with the fine decoder, the model results in small floats and noisy rendering in the dynamic regions (Figure 9c).
Additionally, we demonstrate experiments where the Gaussian center coordinates are injected into the input of each decoder. As shown in 9d, including the Gaussian coordinates not only degrades the quality of deformation but also reduces the quality of static background areas, supporting our argument that coordinate dependency should be removed from the deformation function.
5 Conclusion and Limitation
We propose a per-Gaussian deformation for 3DGS that takes per-Gaussian embeddings as input instead of using the typical deformation field of deformable 3DGS methods, achieving high performance. We enhance the quality by decomposing the dynamic changes into coarse and fine deformation. However, our method also has limitations. The dynamic regions of our rendering results tend to be blurred when there is significant motion between frames, as in the other baselines (Figure 10). It is expected to be resolved with additional supervision like correspondences in rapidly changing dynamic regions. In addition, our method tends to be slower in rendering compared to existing Gaussian Splatting methods. This could be improved by skipping the deformation prediction for static areas and pruning unnecessary points after training, as mentioned before.
Acknowledgments
This work is supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. 2017-0-00072, Development of Audio/Video Coding and Light Field Media Fundamental Technologies for Ultra Realistic Tera-media)
Appendix 0.A More Analysis
0.A.1 Effect of multi-view DSSIM loss
As shown in Figure S1, we observe that DSSIM loss helps to learn difficult areas in the scene, such as the details of the background. However, optimizing DSSIM loss at every step [8] is computationally expensive, and the introduction of DSSIM loss increases the number of Gaussians, further extending the training time. Our multi-view DSSIM loss achieves a similar level of quality to what is computed at every step while reducing the number of computations for DSSIM.
0.A.2 Effect of embedding size
| dim/ dim | PSNR | SSIM | LPIPS | dim/ dim | PSNR | SSIM | LPIPS |
| 256 / 16 | 32.79 | 0.960 | 0.030 | 128 / 32 | 33.23 | 0.962 | 0.028 |
| 256 / 32 | 33.28 | 0.963 | 0.028 | 256 / 32 | 33.28 | 0.963 | 0.028 |
| 256 / 64 | 32.60 | 0.960 | 0.031 | 512 / 32 | 32.80 | 0.961 | 0.031 |
We report the performance variation by changing the size of per-Gaussian and temporal embeddings in Table S1. We observe that the model is sensitive to the size of while it is less sensitive to the size of . We choose the dimensions of and to be 32 and 256, respectively.
Appendix 0.B Training Details
0.B.1 Details of Efficient Training Strategy
For camera sampling, we exclude cameras close to the previous camera center for the next sampling candidates. This is achieved by calculating the pairwise distance between all camera origin and excluding the cameras below the 40 percentile of distances.
For frame sampling, we record the L1 error between rendered images and ground truth for each frame during training. We treat this as a probability distribution for frame sampling to focuse on frames with higher errors. For the first 10K iterations, we randomly sample the frames, and then alternately use random sampling and error-based frame sampling.
Finally, we use multi-view DSSIM loss with L1 loss to induce densification in areas where the model struggles, such as background or dynamic regions. Due to the computational cost of SSIM, we fix the frame obtained by loss-based sampling for every 50 iterations, and minimize the multi-view DSSIM loss through camera sampling during the next 5 iterations. Similar to frame sampling, DSSIM loss is applied after the first 10K iterations.
0.B.2 Implementation Details
In all experiments, we use a 32-dimensional vector for Gaussian embeddings and a 256-dimensional vector for temporal embeddings . The decoder MLP consists 128 hidden units and the MLP head for Gaussian parameters are both composed of 2 layers with 128 hidden units. For Technicolor and HyperNeRF datasets, we use a 1-layer decoder MLP. To efficiently and stabilize the initial training, we start the temporal embedding grid at the same time resolution as and gradually increase it to resolution over 10K iterations, with set to 150 for 300 frames. The learning rate of our deformation decoder starts at and exponentially decays to over the total training iterations. The learning rate of the temporal embedding follows that of the deformation decoder, while the learning rate of per-Gaussian embedding is set to . We eliminate the opacity reset step of 3DGS and instead add a loss that minimizes the mean of Gaussian opacities with a weight of . We empirically set the start of the training strategy to 10K iterations. Following Sync-NeRF, we introduce a learnable time offset to each camera in Neural 3D Video and Technicolor Light Field dataset.
For initialization points, we downsample point clouds obtained from COLMAP dense scene reconstruction. For Neural 3D Video dataset, we train for 80K iterations; for Technicolor dataset, 100K for the Birthday and Painter scenes and 120K for the rest; for the HyperNeRF dataset, 80K for the banana scene and 60K for the others. We periodically use the DSSIM loss at specific steps, excluding the HyperNeRF dataset. For calculating computational cost, the Neural 3D Video dataset was measured using A6000, while the Technicolor Light Field Dataset and the HyperNeRF Dataset were measured using 3090.
Appendix 0.C More Results
Table S2-S4 report the quantitative results of each scene on Neural 3D Video Dataset, Technicolor Light Field Dataset, and HyperNeRF Dataset. In most scenes, our model shows better perceptual performance than the baselines. Additionally, we show the entire rendering results in Figure S2-S4.
0.C.1 Per-Scene Results of Neural 3D Video Dataset
| Average | coffee_martini | cook_spinach | cut_roasted_beef | flame_salmon_1 | ||||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
| MixVoxels | 30.30 | 0.918 | 29.44 | 0.916 | 29.97 | 0.934 | 32.58 | 0.938 | 30.50 | 0.918 |
| K-Planes | 30.86 | 0.939 | 29.66 | 0.926 | 31.82 | 0.943 | 31.82 | 0.966 | 30.68 | 0.928 |
| HyperReel | 30.37 | 0.921 | 28.37 | 0.892 | 32.30 | 0.941 | 32.92 | 0.945 | 28.26 | 0.882 |
| 4DGS | 30.71 | 0.935 | 28.44 | 0.919 | 33.10 | 0.953 | 33.32 | 0.954 | 28.80 | 0.906 |
| 4DGaussians | 31.19 | 0.940 | 28.63 | 0.918 | 33.54 | 0.956 | 34.18 | 0.959 | 29.25 | 0.929 |
| Ours | 31.42 | 0.945 | 29.33 | 0.931 | 33.19 | 0.957 | 33.25 | 0.957 | 29.72 | 0.935 |
| flame_salmon_2 | flame_salmon_3 | flame_salmon_4 | flame_steak | sear_steak | ||||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
| MixVoxels | 30.53 | 0.915 | 27.83 | 0.853 | 29.49 | 0.899 | 30.74 | 0.945 | 31.61 | 0.949 |
| K-Planes | 29.98 | 0.924 | 30.10 | 0.924 | 30.37 | 0.922 | 31.85 | 0.969 | 31.48 | 0.951 |
| HyperReel | 28.80 | 0.911 | 28.97 | 0.911 | 28.92 | 0.908 | 32.20 | 0.949 | 32.57 | 0.952 |
| 4DGS | 28.21 | 0.913 | 28.68 | 0.917 | 28.45 | 0.913 | 33.55 | 0.961 | 34.02 | 0.963 |
| 4DGaussians | 29.28 | 0.929 | 29.40 | 0.927 | 29.13 | 0.925 | 33.90 | 0.961 | 33.37 | 0.960 |
| Ours | 29.76 | 0.931 | 30.12 | 0.935 | 30.28 | 0.934 | 33.55 | 0.963 | 33.55 | 0.963 |
0.C.2 Per-Scene Results of Technicolor Light Field Dataset
| Average | Birthday | Fabien | ||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
| DyNeRF | 31.80 | - | 29.20 | - | 32.76 | - |
| HyperReel | 32.32 | 0.899 | 30.57 | 0.918 | 32.49 | 0.863 |
| 4DGaussians | 29.62 | 0.840 | 28.03 | 0.862 | 33.36 | 0.865 |
| Ours | 33.38 | 0.907 | 33.34 | 0.951 | 34.45 | 0.875 |
| Painter | Theater | Train | ||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
| DyNeRF | 35.95 | - | 29.53 | - | 31.58 | - |
| HyperReel | 35.51 | 0.924 | 33.76 | 0.897 | 29.30 | 0.894 |
| 4DGaussians | 34.52 | 0.899 | 28.67 | 0.835 | 23.54 | 0.756 |
| Ours | 36.63 | 0.923 | 30.64 | 0.864 | 31.84 | 0.922 |
0.C.3 Per-Scene Results of HyperNeRF Dataset
| Average | Broom | 3D Printer | Chicken | Banana | ||||||
| PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | PSNR | SSIM | |
| Nerfies | 22.23 | - | 19.3 | - | 20.00 | - | 26.90 | - | 23.30 | - |
| HyperNeRF DS | 22.29 | 0.598 | 19.51 | 0.210 | 20.04 | 0.635 | 27.46 | 0.828 | 22.15 | 0.719 |
| TiNeuVox | 24.20 | 0.616 | 21.28 | 0.307 | 22.80 | 0.725 | 28.22 | 0.785 | 24.50 | 0.646 |
| D3DGS | 22.40 | 0.598 | 20.48 | 0.313 | 20.38 | 0.644 | 22.64 | 0.601 | 26.10 | 0.832 |
| 4DGaussians | 25.03 | 0.682 | 22.01 | 0.367 | 21.99 | 0.705 | 28.47 | 0.806 | 27.28 | 0.845 |
| Ours | 25.53 | 0.694 | 21.97 | 0.371 | 22.36 | 0.709 | 29.02 | 0.831 | 28.77 | 0.866 |

### v2

@@
 Abstract
-As 3D Gaussian Splatting (3DGS) provides fast and high-quality novel view synthesis, it is a natural extension to deform a canonical 3DGS to multiple frames.
-However, previous works fail to accurately reconstruct dynamic scenes, especially 1) static parts moving along nearby dynamic parts, and 2) some dynamic areas are blurry.
-We attribute the failure to the wrong design of the deformation field, which is built as a coordinate-based function.
+As 3D Gaussian Splatting (3DGS) provides fast and high-quality novel view synthesis, it is a natural extension to deform a canonical 3DGS to multiple frames for representing a dynamic scene.
+However, previous works fail to accurately reconstruct complex dynamic scenes.
+We attribute the failure to the design of the deformation field, which is built as a coordinate-based function.
 This approach is problematic because 3DGS is a mixture of multiple fields centered at the Gaussians, not just a single coordinate-based framework.
@@
 Moreover, we decompose deformations as coarse and fine deformations to model slow and fast movements, respectively.
-Also, we introduce an efficient training strategy for faster convergence and higher quality.
+Also, we introduce a local smoothness regularization for per-Gaussian embedding to improve the details in dynamic regions.
 Project page: https://jeongminb.github.io/e-d3dgs/
 Keywords:
-Gaussian splatting Dynamic scene reconstruction Novel view synthesis1 Introduction
+Deformable gaussian splatting 4D scene reconstruction Novel view synthesis
+1 Introduction
 Dynamic scene reconstruction from multi-view input videos is an important task in computer vision, as it can be extended to various applications and industries such as mixed reality, content production, etc.
@@
 On the other hand, the recently emerging 3D Gaussian Splatting (3DGS) [10] has the advantage of real-time rendering compared to NeRFs using a differentiable rasterizer for 3D Gaussian primitives.
-3DGS directly optimizes the parameters of 3D Gaussian (position, opacity, anisotropic covariance, and spherical harmonics (SH) coefficients) and renders them via projection and -blending.
+3DGS directly optimizes the parameters of 3D Gaussians (position, opacity, anisotropic covariance, and spherical harmonics coefficients) and renders them via projection and -blending.
 Since 3DGS has the characteristics of continuous volumetric radiance fields, some recent studies[29, 32, 3, 8, 14, 33] represent dynamic scenes by defining a canonical 3DGS and deforming it to individual frames as deformable NeRFs [32] do.
-Specifically, they model the deformation as a function of 4D (x, y, z, t) coordinates with MLPs or grids, i.e., fields, predicting the change in the 3D Gaussian parameters.
-However, it is not appropriate to model the deformation of Gaussian parameters as a field for representing dynamic scenes because the positions of Gaussians change during optimization.
-Furthermore, nearby Gaussians tend to have similar deformations because the deformation function is piecewise continuous.
-As shown in Figure 1, the Gaussians in static regions, e.g., window, wall, and handle, are attracted by the motion of nearby dynamic objects.
-Moreover, the existing deformation-based approaches have an inherent property: all frames share a single canonical set of Gaussians.
-It makes the deformation hard to reconstruct different types or difficulties of movements in the dynamic scenes.
+Specifically, they model the deformation as a function of 4D (x, y, z, t) coordinates with MLPs or grids to predict the change in the 3D Gaussian parameters.
+However, since 3DGS is a mixture of multiple volumetric fields, it is not appropriate to model the deformation of Gaussian parameters with a single coordinate-based network to represent dynamic scenes.
+In addition, existing field-based approaches are constrained by the resolution of the grid which models the deformation field, the capacity of the model, or the frequencies of the input.
+As shown in Figure 1, the existing study does not properly represent complex dynamic scenes, and even introducing an additional feature grid that is twice the maximum resolution has only a slight improvement in performance (See Appendix for more results).
+We alleviate this problem by introducing a novel dynamic representation to deform each Gaussian.
 In this paper, we model the deformation of Gaussians at frames as 1) a function of a product space of per-Gaussian embeddings and temporal embeddings.
-We expect this rational design to bring quality improvement by precisely modeling different deformations of different Gaussians without interference due to adjacency.
+We expect this rational design to bring quality improvement by precisely modeling different deformations of different Gaussians.
 Additionally, 2) We decompose temporal variations of the parameters into coarse and fine components, namely coarse-fine deformation.
 The coarse deformation represents large or slow movements in the scene, while fine deformation learns the fast or detailed movements that coarse deformation does not cover.
-Finally, we propose 3) an efficient training strategy that samples difficult parts of the scene more frequently and facilitates the densification of Gaussians.
-In our experiments, we observe that our per-Gaussian embeddings and coarse-fine deformation improve the deformation quality, while an efficient training strategy enables faster convergence.
-Additionally, our approach outperforms baselines in representing dynamic scenes, capturing fine details in dynamic regions, and excelling even under challenging camera settings.
+Finally, we propose 3) a local smoothness regularization for per-Gaussian embedding to ensure the deformations of neighboring Gaussians are similar.
+In our experiments, we observe that our per-Gaussian embeddings, coarse-fine deformation, and regularization improve the deformation quality.
+Our approach outperforms baselines in capturing fine details in dynamic regions and excels even under challenging camera settings.
+Additionally, our method also achieves fast rendering speed and relatively low capacity.
 2 Related Work
-In this section, we review methods for dynamic scene reconstruction that deforms 3D canonical space and methods for reconstructing dynamic scenes utilizing dynamic 3D Gaussians.
-Afterward, we review the methods using embedding and efficient sampling techniques.
+In this section, we review methods for dynamic scene reconstruction that deform 3D canonical space and methods for reconstructing dynamic scenes utilizing dynamic 3D Gaussians.
+Afterward, we review methods that use embeddings and spatial relationships of Gaussians.
 2.0.1 Deforming 3D Canonical Space
@@
 4DGaussians [29] and D3DGS [32] reconstruct the dynamic scene with a deformation network which inputs the center position of the canonical 3D Gaussians and timestamps.
-In contrast, we demonstrate a novel deformation scheme to represent the deformation as a function of a product space of per-Gaussian latent embeddings and temporal embeddings.
+MoDGS[15] learns the mapping between canonical space and space at a specific timestamp through invertible MLP.
+In contrast, we demonstrate a novel deformation representation as a function of a product space of per-Gaussian latent embeddings and temporal embeddings.
 2.0.2 Dynamic 3D Gaussians
 To extend the fast rendering speed of 3D Gaussian Splatting[10] into dynamic scene reconstructions.
-4DGaussians[25] decodes features from multi-resolution HexPlanes for temporal deformation of 3D Gaussians.
+4DGaussians[29] decodes features from multi-resolution HexPlanes [2] for temporal deformation of 3D Gaussians.
 While D3DGS[32] uses an implicit function that processes the time and location of the Gaussian.
-4DGS[26] decomposed the 4D Gaussians into a time-conditioned 3D Gaussians and a marginal 1D Gaussians.
+4DGS[30] decomposes the 4D Gaussians into a time-conditioned 3D Gaussians and a marginal 1D Gaussians.
 STG[13] represents changes in 3D Gaussian over time through a temporal opacity and a polynomial function for each Gaussian.
@@
 Some studies incorporate latent embeddings to represent different states of the static and dynamic scene.
-NeRF-W[13] and Block-NeRF[22] employ per-image appearance embeddings to capture different states of a scene, representing the dynamic scenes from in-the-wild image collections.
+NeRF-W[18] and Block-NeRF[25] employ per-image embeddings to capture different appearances of a scene, representing the scenes from in-the-wild image collections.
 DyNeRF and MixVoxels[12, 27] employ a temporal embedding for each frame to represent dynamic scenes.
@@
 Sync-NeRF[11] introduces time offset to calibrate the misaligned temporal embeddings on dynamic scenes from unsynchronized videos.
-We introduce per-Gaussian latent embedding to encode the changes over time of each Gaussian, and use temporal embeddings to represent different states in each frame of the scene.
-2.0.4 Efficient Training Strategy
-To accelerate convergence and enhance performance, ActiveNeRF[15] selects the view by reducing the uncertainty of the input views.
-DS-NeRF and TermiNeRF[2, 18] samples the points with the sampling network.
-In the dynamic scene reconstruction, DyNeRF[10] samples rays over time, calculating he weight of each ray based on the residual difference between its color across time.
-STG[11] samples additional Gaussians along the ray within patches exhibiting significant errors.
-On the other hand, we present a simple view sampling method through pairwise camera distances, for the first time, we present a sampling method for a target training frame.
+We introduce per-Gaussian latent embedding to encode the changes over time of each Gaussian and use temporal embeddings to represent different states in each frame of the scene.
+2.0.4 Considering Spatial Relationships of Gaussians
+Scaffold-GS [16] reconstructs 3D scenes by synthesizing Gaussians from anchors, utilizing that the neighboring Gaussians have similar properties.
+SAGS [26] creates a graph based on k-nearest neighbors (KNN) so that each Gaussian is optimized while considering its neighboring Gaussians.
+In dynamic scene reconstruction, SC-GS [9] and GaussianPrediction [35] deform Gaussians by combining the deformations of key point Gaussians.
+Dynamic 3D Gaussians [17] utilizes regularization to encourage that Gaussians and their neighboring Gaussians deform with local rigidity.
+Similarly, we propose a local smoothness regularization that encourages neighboring Gaussians to have similar embeddings, resulting in similar deformations.
 3 Method
@@
 Next, we introduce our overall framework, embedding-based deformation for Gaussians (Section 3.2) and coarse-fine deformation scheme consisting of coarse and fine deformation functions (Section 3.3).
-Finally, we present an efficient training strategy for faster convergence and higher performance (Section 3.4).
+Finally, we present a local smoothness regularization for per-Gaussian embeddings to achieve better details on dynamic regions (Section 3.4).
 3.1 Preliminary: 3D Gaussian Splatting
@@
 By its efficient rasterization, the optimized model enables real-time rendering of high-quality images.
-Each 3D Gaussian function at the point consist with position , rotation , and scale :.
+Each 3D Gaussian kernel at the point consist with position , rotation , and scale :
 To projecting 3D Gaussians to 2D for rendering, covariance matrix are calculated by viewing transform and the Jacobian of the affine approximation of the projective transfomation[36] as follows:
-Blending depth-ordered projected points that overlap the pixel, the Gaussian function is multiplied by the opacity of the Gaussian and calculates the pixel color with the color of the Gaussian :
-The color of Gaussian is determined using the SH coefficient with view direction.
+Blending depth-ordered projected points that overlap the pixel, the Gaussian kernel is multiplied by the opacity of the Gaussian and calculates the pixel color with the color of the Gaussian :
+The color of Gaussian is determined using the SH coefficient with the viewing direction.
 3.2 Embedding-Based Deformation for Gaussians
@@
 As a result, the Gaussian parameters at frame are determined by adding to the canonical Gaussian parameters (Figure 2c).
-We jointly optimize the per-Gaussian embeddings , the deformation function , and the canonical Gaussian parameters to minimize rendering loss.
+We jointly optimize the per-Gaussian embeddings , the deformation function , and the canonical Gaussian parameters to minimize the rendering loss.
 We use the L1 and periodic DSSIM as the rendering loss between the rendered image and the ground truth image.
@@
 Refer to the Ablation study section for more details.
-3.4 Efficient Training Strategy
-Last but not least, we introduce a training strategy for faster convergence and higher performance.
-The first is to evenly cover multi-view camera perspectives and exclude the camera index that was used in the previous iteration.
-We pre-compute pairwise distances of all camera origins before the start of the training.
-Then we exclude the cameras with distance to the previous camera less than 40 percentile of all distances.
-The second is to sample frames within the target cameras.
-Different frames may have different difficulties for reconstruction.
-We measure the difficulty of a frame by the magnitude of error at the frame.
-Then, we sample training frames with a categorical distribution which is proportional to the magnitude of error.
-For multi-view videos, we average the errors from all viewpoints.
-We use this sampling every other iteration with random sampling.
-The third is to change the policy for Gaussian densification.
-Previous methods invoke the densification by minimizing L1 loss [25] or L1 loss and DSSIM loss for every iteration [27, 28, 11].
-We observe that the DSSIM loss improves the visual quality in the background but takes longer training time.
-Therefore, we use L1 loss for every iteration and periodically use the additional multi-view DSSIM loss in the frame with high error.
-This encourages densification in the region where the model struggles (See Section 0.A.1).
+3.4 Local Smoothness Regularization
+Neighboring Gaussians constructing dynamic objects tend to exhibit locally similar deformation.
+Inspired by [17], we introduce a local smoothness regularization for per-Gaussian embedding (Figure 2d) to encourage similar deformations between nearby Gaussians and :
+where is the weighting factor and is the Gaussian center.
+We set to 2000 and to 20 following [17].
+To reduce the computational cost, we obtain sets of k-nearest-neighbors only when the densification occurs.
+Note that unlike previous approaches that directly constrain physical properties such as rigidity or rotation, We implicitly induce locally similar deformation by ensuring that per-Gaussian embeddings are locally smooth.
+Our regularization allows better capture of textures and details of dynamic objects.
 4 Experiment
@@
 4.0.1 Baselines
-We choose the start-of-the-art method as a baseline in each datasets.
+We choose the state-of-the-art method as a baseline in each dataset.
 We compared against DyNeRF, NeRFPlayer, MixVoxels, K-Planes, HyperReel, Nerfies, HyperNeRF, and TiNeuVox on the NeRF baseline.
@@
 Also, STG is not a deformable 3D Gaussian approach.
-We followed the official code and configuration, except increasing the training iterations for the Technicolor dataset to 1.5 times that of the 4DGaussians, in comparison to the Neural 3D Video dataset.
+We followed the official code and configuration, except for increasing the training iterations for the Technicolor dataset to 1.5 times that of the 4DGaussians, in comparison to the Neural 3D Video dataset.
 4.0.2 Metrics
@@
 Higher PSNR and SSIM values and lower LPIPS values indicate better visual quality.
-4.1 Effectiveness on dynamic region
-4.1.1 Neural 3D Video Dataset
-[10] includes 20 multi-view videos, with each scene consisting of either 300 frames, except for the flame_salmon scene, which comprises 1200 frames.
-These scenes encompass a relatively long duration and various movements, with some featuring multiple objects in motion.
-We utilized Neural 3D Video dataset to observe the capability to capture dynamic areas.
-Total six scenes (coffee_martini, cook_spinach, cut_roasted_beef, flame_salmon, flame_steak, sear_steak) are evaluated in Figure 4 and Table 1.
-The flame_salmon scene divided into four segments, each containing 300 frames.
-Figure 4 reports the rendering quality.
-Our method successfully reconstructs the fine details in moving areas, outperforming baselines on average metrics across test views.
-Baselines show blurred dynamic areas or severe artifacts in low-light scenes such as cook_spinach and flame_steak.
-4DGS exhibits the disappearance of some static areas.
-In 4DGaussians, a consistent over-smoothing occurs in dynamic areas.
-All baselines experienced reduced quality in reflective or thin dynamic areas like clamps, torches, and windows.
-Table 1 presents quantitative metrics on the average metrics across the test views of all scenes, computational and storage costs on the first fragment of flame_salmon scene.
-Refer to the Appendix for per-scene details.
-Our method demonstrates superior reconstruction quality across all metrics compared to baselines.
-As the table shows, NeRF baselines generally required longer training and rendering times.
-While 4DGS shows relatively high reconstruction performance, it demands longer training times and larger VRAM storage capacity compared to other baselines.
-4DGaussians requires lower computational and storage cost but it displays low reconstruction quality in some scenes with rapid dynamics, as shown in the teaser and Figure 4.
 | model | metric | computational cost | ||||
@@
 | 4DGaussians [29] | 30.71 | 0.935 | 0.056 | 50 mins | 51.9 | 59 MB |
-| Ours | 31.42 | 0.945 | 0.037 | 3 hours 35 mins | 19.9 | 137 MB |
+| Ours | 31.31 | 0.945 | 0.037 | 1 hours 52 mins | 74.5 | 35 MB |
+4.1 Effectiveness on Dynamic Region
+4.1.1 Neural 3D Video Dataset
+[12] includes 20 multi-view videos, with each scene consisting of either 300 frames, except for the flame_salmon scene, which comprises 1200 frames.
+These scenes encompass a relatively long duration and various movements, with some featuring multiple objects in motion.
+We utilized the Neural 3D Video dataset to observe the capability to capture dynamic areas.
+Total six scenes (coffee_martini, cook_spinach, cut_roasted_beef, flame_salmon, flame_steak, sear_steak) are evaluated in Figure 3 and Table 1.
+The flame_salmon scene is divided into four segments, each containing 300 frames.
+Table 1 presents quantitative metrics on the average metrics across the test views of all scenes, computational and storage costs on the first fragment of flame_salmon scene.
+Refer to the Appendix for per-scene details.
+Our method demonstrates superior reconstruction quality, FPS, and model size across compared to baselines.
+As the table shows, NeRF baselines generally required longer training and rendering times.
+While 4DGS shows relatively high reconstruction performance, it demands longer training times and larger VRAM storage capacity compared to other baselines.
+4DGaussians requires lower computational and storage costs but it displays low reconstruction quality in some scenes with rapid dynamics, as shown in the teaser and Figure 3.
+Figure 3 reports the rendering quality.
+Our method successfully reconstructs the fine details in moving areas, outperforming baselines on average metrics across test views.
+Baselines show blurred dynamic areas or severe artifacts in low-light scenes such as cook_spinach and flame_steak.
+4DGS exhibits the disappearance of some static areas.
+In 4DGaussians, a consistent over-smoothing occurs in dynamic areas.
+All baselines experienced reduced quality in reflective or thin dynamic areas like clamps, torches, and windows.
 4.1.2 Technicolor Light Field Dataset
-[20] is a multi-view dataset captured with a time-synchronized camera rig, containing intricate details.
-We train ours and the baselines on 50 frames of five commonly used scenes (Birthday, Fabien, Painter, Theater, Trains) using full-resolution videos at pixels, with the second row and second column cameras used as test views.
-As shown in Figure 5, HyperReel produces noisy artifacts due to incorrect predictions of the displacement vector.
-4DGaussians fails to capture fine details in dynamic areas, exhibiting over-smoothing results.
-All baselines struggle to accurately reconstruct rapidly moving thin areas like fingers.
-Table 2 reports the average metrics across the test views of all scenes, computational and storage costs on Painter scene.
-HyperReel demonstrates overall high-quality results but struggles with relatively slow training times and FPS, and a larger model size.
-4DGaussians exhibits fast training times and FPS but significantly underperforms in reconstructing fine details compared to other baselines.
-However, our method demonstrates superior reconstruction quality and faster FPS compared to the baselines.
 | model | metric | computational cost | ||||
@@
 | 4DGaussians | 29.62 | 0.844 | 0.176 | 25 mins | 34.8 | 51 MB |
-| Ours | 33.38 | 0.907 | 0.103 | 1 hours 30 mins | 36.5 | 56 MB |
+| Ours | 33.24 | 0.907 | 0.100 | 2 hours 55 mins | 60.8 | 77 MB |
+[23] is a multi-view dataset captured with a time-synchronized camera rig, containing intricate details.
+We train ours and the baselines on 50 frames of five commonly used scenes (Birthday, Fabien, Painter, Theater, Trains) using full-resolution videos at pixels, with the second row and second column cameras used as test views.
+Table 2 reports the average metrics across the test views of all scenes, computational and storage costs on Painter scene.
+HyperReel demonstrates overall high-quality results but struggles with relatively slow training times and FPS, and a larger model size.
+4DGaussians exhibits fast training times and FPS but significantly underperforms in reconstructing fine details compared to other baselines.
+However, our method demonstrates superior reconstruction quality and faster FPS compared to the baselines.
+As shown in Figure 4, HyperReel produces noisy artifacts due to incorrect predictions of the displacement vector.
+4DGaussians fails to capture fine details in dynamic areas, exhibiting over-smoothing results.
+All baselines struggle to accurately reconstruct rapidly moving thin areas like fingers.
 4.2 Scenes with Challenging Camera Setting
-4.2.1 HyperNeRF dataset
-includes videos using two Pixel 3 phones rigidly mounted on a handheld capture rig.
-We train on all frames of four scenes (3D Printer, Banana, Broom, Chicken) at a resolution downsampled by a factor of two to .
-Due to memory constraints, D3DGS is trained on images further downsampled to a quarter of their original size.
-Figure 6 shows that the previous methods struggle to reconstruct fast-moving parts such as fingers and broom.
-Especially D3DGS deteriorates in Broom.
-Table 3 reports the average metrics across the test views of all scenes, computational and storage costs on Broom scene.
-The table shows that our method outperforms the reconstruction performance with previous methods along with compact model size and faster FPS.
 | model | metric | computational cost | ||||
@@
 | 4DGaussians | 25.03 | 0.682 | 0.281 | 50 mins | 96.3 | 60 MB |
-| Ours | 25.53 | 0.694 | 0.242 | 1 hours 20 mins | 30.9 | 34 MB |
+| Ours | 25.43 | 0.697 | 0.231 | 1 hours 15 mins | 139.3 | 33 MB |
+4.2.1 HyperNeRF Dataset
+includes videos captured using two phones rigidly mounted on a handheld rig.
+We train on all frames of four scenes (3D Printer, Banana, Broom, Chicken) at a resolution downsampled by half to .
+Due to memory constraints, D3DGS is trained on images downsampled by a quarter.
+The table shows that our method outperforms the reconstruction performance with previous methods along with compact model size and faster FPS.
+Figure 5 shows that previous methods struggle to reconstruct fast-moving parts such as fingers and broom.
+Especially D3DGS deteriorates in Broom scene.
+Table 3 reports the average metrics across the test views of all scenes, computational and storage costs on Broom scene.
 4.3 Analyses and Ablation Study
-4.3.1 Improving Rendering Speed
-Our method can enhance rendering speed by omitting the deformation of Gaussians in static areas, which only require a single computation of parameter changes from the canonical space.
-To identify static Gaussians, we sum up the L1 norms of each Gaussian parameter over time and select Gaussians below the percentile based on the aggregated value.
-As Table 4 reports, we can obtain almost identical rendering quality to the original, despite omitting decoder evaluations for 90% of the total Gaussian.
-The experiments are conducted on flame_salmon_1 scene with a single NVIDIA RTX 4090.
-| Skip deformation | PSNR | SSIM | LPIPS | FPS |
-| 0% (original) | 29.72 | 0.935 | 0.038 | 17.1 |
-| 50% | 29.72 | 0.935 | 0.038 | 25.7 |
-| 70% | 29.72 | 0.935 | 0.038 | 26.0 |
-| 90% | 29.68 | 0.935 | 0.038 | 28.4 |
 4.3.1 Deformation components
@@
 | Method | PSNR | SSIM | LPIPS |
-| Ours Full | 29.08 | 0.924 | 0.056 |
-| w/o Coarse-Fine deformation | 28.74 | 0.922 | 0.060 |
-| w/o Efficient training strategy | 28.89 | 0.920 | 0.056 |
+| Ours | 29.70 | 0.933 | 0.041 |
+| coarse deformation only | 29.48 | 0.931 | 0.044 |
+| fine deformation only | 29.23 | 0.932 | 0.043 |
+| Ours + injection | 29.60 | 0.931 | 0.045 |
+4.3.2 Ablation study
+We report the results of an ablation study on the deformation decoder in Figure 8 and Table 4.
+First, our full method (using both coarse and fine decoders) produces clear rendering results and models dynamic variations well (Figure 8a).
+Training only with the coarse or fine decoder leads to blurred dynamic areas and a failure to accurately capture detailed motion (Figure 8b-c).
+Additionally, we demonstrate experiments where the Gaussian center coordinates are injected into the input of each decoder.
+As shown in Figure 8d, including the Gaussian coordinates degrades the quality of deformation, supporting our argument that coordinate dependency should be removed from the deformation function.
 | Method | PSNR | SSIM | LPIPS |
-| Ours Full | 30.28 | 0.934 | 0.047 |
-| coarse deformation only | 29.71 | 0.930 | 0.050 |
-| fine deformation only | 29.62 | 0.929 | 0.048 |
-| Ours + injection | 28.17 | 0.916 | 0.057 |
-4.3.3 Ablation study
-To verify the effectiveness of each component of our method, we report the performance when each component is removed.
-To observe the impact of the coarse-fine deformation, we train a total of 100K iterations on flame_salmon_4 scene.
-To show the efficiency of our training strategy, we evaluate the rendering results from checkpoints of 30K iterations.
-As shown in Table 5, we confirm that the performance decreases when each component is removed.
-Furthermore, we report the results of an ablation study on the deformation decoder in Figure 9 and Table 6.
-First, our full method (using both coarse and fine decoders) produces clear rendering results and models dynamic variations well (Figure 9a).
-Training only with the coarse decoder leads to blurred dynamic areas and a failure to accurately capture detailed motion (Figure 9)b.
-On the other hand, only with the fine decoder, the model results in small floats and noisy rendering in the dynamic regions (Figure 9c).
-Additionally, we demonstrate experiments where the Gaussian center coordinates are injected into the input of each decoder.
-As shown in 9d, including the Gaussian coordinates not only degrades the quality of deformation but also reduces the quality of static background areas, supporting our argument that coordinate dependency should be removed from the deformation function.
+| Ours w/o embedding reg | 32.26 | 0.951 | 0.037 |
+| + our embedding reg | 32.34 | 0.952 | 0.036 |
+| + physically-based reg | 32.08 | 0.950 | 0.036 |
+Furthermore, we report the results of an ablation study on the local smoothness regularization for per-Gaussian embeddings.
+As shown in Figure 9, our regularization improves the details and texture quality of moving objects.
+In Table 5, we show a performance comparison between the proposed regularization and the existing physically-based regularizations.
+To apply the method proposed by Luiten et al.[17] to ours, we make some modifications: 1) Like our method, we find the set of k-nearest neighbors only when the densification occurs to reduce the computational cost.
+2) For long-term local-isometry loss, we use the time of the video frame used in the previous training step instead of using the time of the first frame.
+Our regularization is simple and shows better performance compared to the previous method.
 5 Conclusion and Limitation
-We propose a per-Gaussian deformation for 3DGS that takes per-Gaussian embeddings as input instead of using the typical deformation field of deformable 3DGS methods, achieving high performance.
-We enhance the quality by decomposing the dynamic changes into coarse and fine deformation.
-However, our method also has limitations.
-The dynamic regions of our rendering results tend to be blurred when there is significant motion between frames, as in the other baselines (Figure 10).
-It is expected to be resolved with additional supervision like correspondences in rapidly changing dynamic regions.
-In addition, our method tends to be slower in rendering compared to existing Gaussian Splatting methods.
-This could be improved by skipping the deformation prediction for static areas and pruning unnecessary points after training, as mentioned before.
-Acknowledgments
+We propose a per-Gaussian deformation for 3DGS that takes per-Gaussian embeddings as input, instead of using the typical deformation fields from previous deformable 3DGS works, resulting in high performance.
+We enhance the reconstruction quality by decomposing the dynamic changes into coarse and fine deformation.
+However, our method learns inappropriate Gaussian deformation with casually captured monocular videos[7], like other baselines.
+We plan to address it in future work by introducing useful prior for monocular video settings.
+Acknowledgements
 This work is supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.
 2017-0-00072, Development of Audio/Video Coding and Light Field Media Fundamental Technologies for Ultra Realistic Tera-media)
-Appendix 0.A More Analysis
-0.A.1 Effect of multi-view DSSIM loss
-As shown in Figure S1, we observe that DSSIM loss helps to learn difficult areas in the scene, such as the details of the background.
-However, optimizing DSSIM loss at every step [8] is computationally expensive, and the introduction of DSSIM loss increases the number of Gaussians, further extending the training time.
-Our multi-view DSSIM loss achieves a similar level of quality to what is computed at every step while reducing the number of computations for DSSIM.
-0.A.2 Effect of embedding size
+Appendix 0.A Training Details
+0.A.1 Efficient Training Strategy
+We introduce a training strategy for faster convergence and higher performance.
+The first is to evenly cover multi-view camera perspectives and exclude the camera index that was used in the previous iteration.
+We pre-compute pairwise distances of all camera origins before the start of the training.
+Then we exclude the cameras with distance to the previous camera less than 40 percentile of all distances.
+The second is to sample frames within the target cameras.
+Different frames may have different difficulties for reconstruction.
+We measure the difficulty of a frame by the magnitude of error at the frame.
+Then, we sample training frames with a categorical distribution which is proportional to the magnitude of error.
+For multi-view videos, we average the errors from all viewpoints.
+For the first 10K iterations, we randomly sample the frames, and then alternately use random sampling and error-based frame sampling.
+The third is to change the policy for Gaussian densification.
+Previous methods invoke the densification by minimizing L1 loss [29] or L1 loss and DSSIM loss for every iteration [31, 32, 13].
+We observe that the DSSIM loss improves the visual quality in the background but takes longer training time.
+Therefore, we use L1 loss for every iteration and periodically use the additional multi-view DSSIM loss in the frame with high error.
+This encourages densification in the region where the model struggles.
+We fix the frame obtained by loss-based sampling for every 50 iterations, and minimize the multi-view DSSIM loss through camera sampling during the next 5 iterations.
+Similar to frame sampling, DSSIM loss is applied after the first 10K iterations.
+0.A.2 Implementation Details
 | dim/ dim | PSNR | SSIM | LPIPS | dim/ dim | PSNR | SSIM | LPIPS |
-| 256 / 16 | 32.79 | 0.960 | 0.030 | 128 / 32 | 33.23 | 0.962 | 0.028 |
-| 256 / 32 | 33.28 | 0.963 | 0.028 | 256 / 32 | 33.28 | 0.963 | 0.028 |
-| 256 / 64 | 32.60 | 0.960 | 0.031 | 512 / 32 | 32.80 | 0.961 | 0.031 |
+| 256 / 16 | 33.53 | 0.962 | 0.030 | 128 / 32 | 33.61 | 0.964 | 0.029 |
+| 256 / 32 | 33.57 | 0.964 | 0.028 | 256 / 32 | 33.57 | 0.964 | 0.028 |
+| 256 / 64 | 33.71 | 0.964 | 0.029 | 512 / 32 | 33.55 | 0.964 | 0.028 |
 We report the performance variation by changing the size of per-Gaussian and temporal embeddings in Table S1.
-We observe that the model is sensitive to the size of while it is less sensitive to the size of .
-We choose the dimensions of and to be 32 and 256, respectively.
-Appendix 0.B Training Details
-0.B.1 Details of Efficient Training Strategy
-For camera sampling, we exclude cameras close to the previous camera center for the next sampling candidates.
-This is achieved by calculating the pairwise distance between all camera origin and excluding the cameras below the 40 percentile of distances.
-For frame sampling, we record the L1 error between rendered images and ground truth for each frame during training.
-We treat this as a probability distribution for frame sampling to focuse on frames with higher errors.
-For the first 10K iterations, we randomly sample the frames, and then alternately use random sampling and error-based frame sampling.
-Finally, we use multi-view DSSIM loss with L1 loss to induce densification in areas where the model struggles, such as background or dynamic regions.
-Due to the computational cost of SSIM, we fix the frame obtained by loss-based sampling for every 50 iterations, and minimize the multi-view DSSIM loss through camera sampling during the next 5 iterations.
-Similar to frame sampling, DSSIM loss is applied after the first 10K iterations.
-0.B.2 Implementation Details
-In all experiments, we use a 32-dimensional vector for Gaussian embeddings and a 256-dimensional vector for temporal embeddings .
+The embedding size does not significantly affect the overall performance.
+In all experiments, considering the capacity and details, we use a 32-dimensional vector for Gaussian embeddings and a 256-dimensional vector for temporal embeddings .
 The decoder MLP consists 128 hidden units and the MLP head for Gaussian parameters are both composed of 2 layers with 128 hidden units.
@@
 We empirically set the start of the training strategy to 10K iterations.
-Following Sync-NeRF, we introduce a learnable time offset to each camera in Neural 3D Video and Technicolor Light Field dataset.
+For the Neural 3D Video dataset, we follow Sync-NeRF[11] to introduce trainable time offsets to each camera.
+In addition, we learn the deformation of color, scale, and opacity from 5K iterations and perform densification and pruning.
 For initialization points, we downsample point clouds obtained from COLMAP dense scene reconstruction.
-For Neural 3D Video dataset, we train for 80K iterations; for Technicolor dataset, 100K for the Birthday and Painter scenes and 120K for the rest; for the HyperNeRF dataset, 80K for the banana scene and 60K for the others.
+For the Neural 3D Video dataset, we train for 80K iterations; for the Technicolor dataset, 100K for the Birthday and Painter scenes and 120K for the rest; for the HyperNeRF dataset, 80K for the Banana scene and 60K for the others.
 We periodically use the DSSIM loss at specific steps, excluding the HyperNeRF dataset.
 For calculating computational cost, the Neural 3D Video dataset was measured using A6000, while the Technicolor Light Field dataset and the HyperNeRF dataset were measured using 3090.
-Appendix 0.C More Results
-Table S2-S4 report the quantitative results of each scene on Neural 3D Video Dataset, Technicolor Light Field Dataset, and HyperNeRF Dataset.
+For local smoothness embedding regularization, we set the weight of loss to 1 on the Neural 3D Video dataset and 0.1 on the Technicolor Light Field dataset and the HyperNeRF dataset.
+Appendix 0.B More Results
+We show the additional rendering of 4DGasussians [29] in Figure S2.
+And, we show the entire rendering results of ours in Figure S3-S5.
 In most scenes, our model shows better perceptual performance than the baselines.
-Additionally, we show the entire rendering results in Figure S2-S4.
-0.C.1 Per-Scene Results of Neural 3D Video Dataset
+Table S2-S4 report the quantitative results of each scene on the Neural 3D Video dataset, Technicolor Light Field dataset, and HyperNeRF dataset.
+0.B.1 Additional Results of 4DGaussians
+The results of 4DGaussians are produced using the official repository.
+In the teaser, we choose the third fragment of flame salmon scene which contains more difficult movements than the typical first fragment to demonstrate our superiority.
+0.B.2 Per-Scene Results of Neural 3D Video Dataset
 | Average | coffee_martini | cook_spinach | cut_roasted_beef | flame_salmon_1 | ||||||
@@
 | HyperReel | 30.37 | 0.921 | 28.37 | 0.892 | 32.30 | 0.941 | 32.92 | 0.945 | 28.26 | 0.882 |
-| 4DGS | 30.71 | 0.935 | 28.44 | 0.919 | 33.10 | 0.953 | 33.32 | 0.954 | 28.80 | 0.906 |
-| 4DGaussians | 31.19 | 0.940 | 28.63 | 0.918 | 33.54 | 0.956 | 34.18 | 0.959 | 29.25 | 0.929 |
-| Ours | 31.42 | 0.945 | 29.33 | 0.931 | 33.19 | 0.957 | 33.25 | 0.957 | 29.72 | 0.935 |
+| 4DGaussians | 30.71 | 0.935 | 28.44 | 0.919 | 33.10 | 0.953 | 33.32 | 0.954 | 28.80 | 0.906 |
+| 4DGS | 31.19 | 0.940 | 28.63 | 0.918 | 33.54 | 0.956 | 34.18 | 0.959 | 29.25 | 0.929 |
+| Ours | 31.31 | 0.945 | 29.10 | 0.931 | 32.96 | 0.956 | 33.57 | 0.958 | 29.61 | 0.936 |
 | flame_salmon_2 | flame_salmon_3 | flame_salmon_4 | flame_steak | sear_steak | ||||||
@@
 | HyperReel | 28.80 | 0.911 | 28.97 | 0.911 | 28.92 | 0.908 | 32.20 | 0.949 | 32.57 | 0.952 |
-| 4DGS | 28.21 | 0.913 | 28.68 | 0.917 | 28.45 | 0.913 | 33.55 | 0.961 | 34.02 | 0.963 |
-| 4DGaussians | 29.28 | 0.929 | 29.40 | 0.927 | 29.13 | 0.925 | 33.90 | 0.961 | 33.37 | 0.960 |
-| Ours | 29.76 | 0.931 | 30.12 | 0.935 | 30.28 | 0.934 | 33.55 | 0.963 | 33.55 | 0.963 |
-0.C.2 Per-Scene Results of Technicolor Light Field Dataset
+| 4DGaussians | 28.21 | 0.913 | 28.68 | 0.917 | 28.45 | 0.913 | 33.55 | 0.961 | 34.02 | 0.963 |
+| 4DGS | 29.28 | 0.929 | 29.40 | 0.927 | 29.13 | 0.925 | 33.90 | 0.961 | 33.37 | 0.960 |
+| Ours | 29.69 | 0.934 | 29.94 | 0.934 | 29.90 | 0.933 | 33.57 | 0.964 | 33.45 | 0.963 |
+0.B.3 Per-Scene Results of Technicolor Light Field Dataset
 | Average | Birthday | Fabien | ||||
@@
 | 4DGaussians | 29.62 | 0.840 | 28.03 | 0.862 | 33.36 | 0.865 |
-| Ours | 33.38 | 0.907 | 33.34 | 0.951 | 34.45 | 0.875 |
+| Ours | 33.23 | 0.907 | 32.90 | 0.951 | 34.71 | 0.885 |
 | Painter | Theater | Train | ||||
@@
 | 4DGaussians | 34.52 | 0.899 | 28.67 | 0.835 | 23.54 | 0.756 |
-| Ours | 36.63 | 0.923 | 30.64 | 0.864 | 31.84 | 0.922 |
-0.C.3 Per-Scene Results of HyperNeRF Dataset
+| Ours | 36.18 | 0.924 | 31.07 | 0.868 | 31.33 | 0.912 |
+0.B.4 Per-Scene Results of HyperNeRF Dataset
 | Average | Broom | 3D Printer | Chicken | Banana | ||||||
@@
 | 4DGaussians | 25.03 | 0.682 | 22.01 | 0.367 | 21.99 | 0.705 | 28.47 | 0.806 | 27.28 | 0.845 |
-| Ours | 25.53 | 0.694 | 21.97 | 0.371 | 22.36 | 0.709 | 29.02 | 0.831 | 28.77 | 0.866 |
+| Ours | 25.43 | 0.697 | 21.84 | 0.371 | 22.34 | 0.715 | 28.75 | 0.836 | 28.80 | 0.867 |